Find 3 consecutive integers such that the sum of the first two is 13 less than the third
Answer by fractalier(6550)

Call the three consecutive integers x, x+1, and x+2.
Thus
x + (x+1) = (x+2)-13
Now solve...
2x+1=x-11
x = -12
Thus the numbers are
-12, -11, and -10.